# Jl. Ganesha 10, Bandung 40132, Indonesia

### v1

@@ 3 Numerical results and discussion
Table 4: Comparison between the hydrostatic Newtonian mass, hydrostatic mass in Rastall gravity for two values of the Rastall parameter, and the lensing mass of galaxy clusters. The Newtonian mass and $r_{500}$ data are taken from Ref. [4], while the lensing mass data are taken from Refs. [57, 58, 59].
Cluster $r_{500}$ $M_{\mathrm{Newt}}$ $M_{hyd}$ $M_{hyd}$ $M_{\mathrm{len}}$
@@
$\lambda=1.49\times 10^{-3}$ $\lambda=1.69\times 10^{-3}$
A383 $944\pm 32$ 3.20 3.70 3.80 5.03
A478 $1337\pm 58$ 7.40 8.57 8.80 8.67
A1795 $1235\pm 36$ 5.80 6.65 6.83 9.20
A1991 $732\pm 33$ 1.40 1.60 1.64 2.50
A2029 $1362\pm 43$ 8.60 9.98 10.24 10.80
In this work, we compute the masses of galaxy clusters at the radius $r_{500}$. As shown in Tables 3 and 4, the values of $r_{500}$ are subject to observational uncertainties. In our analysis, however, the mass is evaluated at the central (best-fit) value of $r_{500}$, while the propagation of the uncertainty in $r_{500}$ is not taken into account. This approach is also employed in Ref. [4].
@@
The formula for the relative likelihood is presented in Appendix A. From Fig. 6(a), it is evident that the standard (non-Rastall) hydrostatic mass yields a slope that deviates significantly from unity, namely $\mathbf{M}=8.24\pm 0.86$. Even when the uncertainty is taken into account, the value $\mathbf{M}=1$ lies well outside the allowed range. Therefore, within the context of the assumption that DM is absent, this case is not viable. In contrast, the assumption of no DM is more compatible with the Rastall gravity framework.
@@
For Rastall gravity with $\lambda=9.55\times 10^{-2}$, the resulting slope is significantly improved compared to the non-Rastall case. Although the value $\mathbf{M}=1$ is still not encompassed by the uncertainty of the slope, it is nevertheless included within the corresponding relative likelihood curve.
@@
In the comparison between the hydrostatic mass and the baryonic mass, the value $\lambda=1.14\times 10^{-1}$ yields the best linear fit, with a slope $\mathbf{M}=1.07\pm 0.11$. The slope is very close to unity, and the value $\mathbf{M}=1$ lies within both the uncertainty range and the corresponding relative likelihood curve. This result indicates that, from a statistical standpoint, Rastall gravity provides a more favorable description of the hydrostatic-baryonic mass relation than standard gravity under the assumption of the absence of DM. Another noteworthy result is that the best-fit slope obtained in this work is closer to unity than those reported in some related studies. In Ref. [4], the best-fit slopes were found to be $0.126\pm 0.086$ for EiBI theory and $0.023\pm 0.119$ for BHG. Meanwhile, Ref. [22] reported a slope of $0.84\pm 0.04$ within the framework of nonlocal gravity.
@@
We also examine whether the hydrostatic mass within the framework of Rastall gravity can help to alleviate the hydrostatic mass bias, an issue that was not resolved in Ref. [16]. To this end, we consider a sample of five galaxy clusters. The values of the Newtonian hydrostatic mass, the Rastall hydrostatic mass, and the corresponding lensing mass for these clusters are presented in Table 4. For this test, we adopt $\lambda=1.49\times 10^{-3}$ and $\lambda=1.69\times 10^{-3}$. Both values lead to an increase in the hydrostatic mass.
@@
The linear fits between the standard (non-Rastall) hydrostatic mass and the observed lensing mass, as well as between the Rastall hydrostatic mass and the observed lensing mass, are shown in Fig. 6(b). The standard non-Rastall hydrostatic mass yields a slope value of $\mathbf{M}=0.85\pm 0.13$. Although this value is already close to unity, we further investigate whether the inclusion of Rastall gravity can improve the linear relation and bring the slope closer to the ideal value $\mathbf{M}=1$, thereby alleviating the hydrostatic mass bias.
@@
In the comparison between the hydrostatic mass and the lensing mass, the value $\lambda=1.69\times 10^{-3}$ yields the best linear fit, with a slope $\mathbf{M}=1.01\pm 0.16$. The slope is very close to unity, even closer to $\mathbf{M}=1$ than the best-fit value obtained from the comparison between the hydrostatic mass and the observed baryonic mass. Moreover, the value $\mathbf{M}=1$ lies within the uncertainty range and is also located very close to the peak of the corresponding relative likelihood curve.
@@
This result indicates that, from a statistical standpoint, Rastall gravity is able to significantly alleviate the hydrostatic mass bias. Nevertheless, a test with a larger sample of galaxy clusters and more lensing mass data is necessary, both to further assess the viability of Rastall gravity and to constrain the value of the Rastall parameter $\lambda$ using observational data. Another noteworthy outcome of this study is that it addresses the issue encountered in Ref. [16], where the hydrostatic mass bias could not be alleviated within the considered framework.
@@ 4 Conclusion
In this study, we have derived the formalism of the hydrostatic mass of galaxy clusters within the framework of Rastall gravity. We find that Rastall gravity has a nontrivial influence on the hydrostatic mass of galaxy clusters. Two scenarios are considered: (i) the absence of dark matter (DM), in which the Rastall hydrostatic mass is compared with the observed baryonic mass; and (ii) the existence of DM, where the Rastall hydrostatic mass is employed to alleviate the hydrostatic mass bias.
@@
On the other hand, using five galaxy clusters as the sample, we find that the best-fit value, $\mathbf{M}=1.01\pm 0.16$, is obtained for $\lambda=1.69\times 10^{-3}$. In this case, unity lies very close to the peak of the corresponding relative likelihood curve in the comparison between the Rastall hydrostatic mass and the observed lensing mass. This result indicates that Rastall gravity can effectively alleviate the hydrostatic mass bias problem. Nevertheless, a larger sample of galaxy clusters in the comparison between the Rastall hydrostatic mass and the gravitational lensing mass is required to further assess the viability of Rastall gravity and to constrain the Rastall parameter $\lambda$ using observational data.
@@
During the preparation of this work, the author used ChatGPT (OpenAI) in order to improve the language and readability of the manuscript. After using this tool, he reviewed and edited the content as needed and take full responsibility for the content of the paper.
@@ Acknowledgements
Appendix A Formula of relative likelihood
$\mathcal{L}_{\mathrm{rel}}=\exp\left[-\frac{1}{2}\left(\frac{\mathbf{M}-\mu}{\sigma}\right)^{2}\right],$ (A1)
Appendix B Statistical data of the linear fits

### v2

@@ 3 Numerical results, statistical evaluations, and discussion
In this work, we compute the masses of galaxy clusters at the radius $r_{500}$. As shown in Tables 4 and 5, the values of $r_{500}$ are subject to observational uncertainties. In our analysis, however, the mass is evaluated at the central (best-fit) value of $r_{500}$, while the propagation of the uncertainty in $r_{500}$ is not taken into account. This approach is also employed in Ref. [4].
Table 5: Comparison between the hydrostatic Newtonian mass, hydrostatic mass in Rastall gravity for two values of the Rastall parameter, and the lensing mass of galaxy clusters. The Newtonian mass and $r_{500}$ data are taken from Ref. [4], while the lensing mass data are taken from Refs. [73, 74, 75, 76].
Cluster $r_{500}$ $M_{\mathrm{Newt}}$ $M_{hyd}^{\text{R}}$ $M_{hyd}^{\text{R}}$ $M_{\mathrm{len}}$
@@
$\lambda=1.14\times 10^{-3}$ $\lambda=1.29\times 10^{-3}$
A133 $1007\pm 41$ 4.10 4.61 4.53 $2.8\pm 1.4$
A383 $944\pm 32$ 3.20 3.55 3.61 5.03 $\pm 1.34$
A478 $1337\pm 58$ 7.40 8.24 8.37 8.67 $\pm 2.01$
A907 $1096\pm 30$ 4.40 4.91 4.99 $3.38\pm 0.94$
A1413 $1299\pm 43$ $9.90$ 11.05 11.23 $7.20\pm 2.00$
A1795 $1235\pm 36$ 5.80 6.39 6.50 9.20 $\pm 2.3$
A1991 $732\pm 33$ 1.40 1.53 1.56 2.50 $\pm 1.75$
A2029 $1362\pm 43$ 8.60 9.59 9.75 10.80 $\pm 2.10$
A2390 $1416\pm 48$ 13.00 14.27 14.51 $10.9\pm 1.7$
@@
Since our aim in the first scenario is to constrain the Rastall parameter within the hydrostatic mass framework under the assumption of no DM, we consider parameter values that yield the best agreement with the observational data. In particular, we focus on two representative values of the Rastall parameter that produce the closest approach to a linear relation with slope near unity. While one of these values provides the overall best-fit, the second value is also retained for comparison purposes, in order to illustrate the sensitivity of the results to variations of the Rastall parameter
To further quantify the consistency between the model and the data, we also evaluate the corresponding relative likelihood. Note that since we are working with Rastall gravity that results in extra degree of freedom brought by Rastall parameter $\lambda$, and the uncertainty always appears in the measurements, we also have to evaluate the goodness-of-fit by calculating the chi-squared $\chi^{2}$ and reduced chi-squared $\chi_{\nu}^{2}$ of every fit.
${\color[rgb]{0,0,0}\mathcal{L}_{\mathrm{rel}}=\exp\left[-\frac{1}{2}\left(\frac{\mathbf{M}-\mu}{\sigma}\right)^{2}\right],}$ (36)
The formula of $\chi^{2}$ is given by [72]
${\color[rgb]{0,0,0}\chi^{2}=\sum^{N}_{n=1}\left(\frac{y_{n}-E_{n}}{\sigma_{n}}\right),}$ (37)
where we have $N$ data values $y_{n}$, measured with uncertainties $\sigma_{n}$, and $E_{n}$ denotes the value from the model. Moreover, $\chi_{\nu}^{2}$ reads
${\color[rgb]{0,0,0}\chi_{\nu}^{2}=\frac{\chi^{2}}{N-\mathscr{P}}.}$ (38)
Here $\mathscr{P}$ denotes the amount of fitted parameters.
@@
From Fig. 1(a), it is evident that the standard (non-Rastall) hydrostatic mass yields a slope that deviates significantly from unity, namely $\mathbf{M}=8.24\pm 0.86$. Even when the uncertainty is taken into account, the value $\mathbf{M}=1$ lies well outside the allowed range. Therefore, the strong mismatch with the observed baryonic mass under the assumption of no DM reflects the well-known discrepancy in which the hydrostatic mass substantially exceeds the baryonic contribution. Moreover, from Fig. 1(b), it is evident that the value $\mathbf{M}=1$ does not lie within the tail of the likelihood distribution, indicating that the standard Newtonian estimate is strongly disfavored by the observed baryonic mass data.
@@
In this calculation, we found that the values $\lambda=9.55\times 10^{-2}$ and $\lambda=1.14\times 10^{-1}$ result in the slopes which are very close to unity, as can be seen in Fig. 1(c) and (e). In Fig. 1(d) and (f), it is obvious that the value $\mathbf{M}=1$ lies within the curve of the likelihood distribution, which means that both values of parameter are favored by the observational baryonic data. For Rastall gravity with $\lambda=9.55\times 10^{-2}$, the resulting slope is significantly improved compared to the non-Rastall case. As we can see in Fig. 1(d), although the value $\mathbf{M}=1$ is still not encompassed by the uncertainty of the slope, it is nevertheless included within the corresponding relative likelihood curve.
@@
In the comparison between the hydrostatic mass and the baryonic mass, the value $\lambda=1.14\times 10^{-1}$ yields the best linear fit, with a slope $\mathbf{M}=1.07\pm 0.11$, as shown by Fig. 1(e). The slope is very close to unity, and the value $\mathbf{M}=1$ lies within both the uncertainty range and the corresponding relative likelihood curve, as shown by Fig. 1(f). This result indicates that, from a statistical standpoint, Rastall gravity provides a more favorable description of the hydrostatic-baryonic mass relation than standard gravity under the assumption of the absence of DM. Another noteworthy result is that the best-fit slope obtained in this work is closer to unity than those reported in some related studies. In Ref. [4], the best-fit slopes were found to be $0.126\pm 0.086$ for EiBI theory and $0.023\pm 0.119$ for BHG. Meanwhile, Ref. [33] reported a slope of $0.84\pm 0.04$ within the framework of nonlocal gravity.
@@
In the second scenario, we also examine whether the hydrostatic mass within the framework of Rastall gravity can help to alleviate the hydrostatic mass bias, an issue that was not resolved in Ref. [27]. To this end, we consider a sample of five galaxy clusters. The values of the Newtonian hydrostatic mass, the Rastall hydrostatic mass, and the corresponding lensing mass for these clusters are presented in Table 5. Note that in Table 5, the measurement uncertainties are treated as symmetric. In the original dataset, some uncertainties are asymmetric. To symmetrize them, we take the mean values of the upper and lower uncertainties. As in the first scenario, the aim at this stage is to constrain the Rastall parameter in the cluster-scale, now under the assumption of the presence of DM, while examining whether the discrepancy between hydrostatic and lensing mass estimates can be alleviated. We consider representative parameter values that yield good agreement with the observational data. For this test, we adopt $\lambda=1.49\times 10^{-3}$ and $\lambda=1.69\times 10^{-3}$. Both values lead to an increase in the hydrostatic mass and result in slopes close to unity.
@@
The linear fits between the standard (non-Rastall) hydrostatic mass and the observed lensing mass, as well as between the Rastall hydrostatic mass and the observed lensing mass, are shown in Fig. 2. The standard non-Rastall hydrostatic mass yields a slope value of $\mathbf{M}=0.88\pm 0.23$, as shown by Fig. 2(a). Although this value is already close to unity, we further investigate whether the inclusion of Rastall gravity can improve the linear relation and bring the slope closer to the ideal value $\mathbf{M}=1$, thereby alleviating the hydrostatic mass bias.
@@
In the comparison between the hydrostatic mass and the lensing mass, the value $\lambda=1.14\times 10^{-3}$ yields the slope $\mathbf{M}=0.98\pm 0.26$, as shown by Fig. 2(c). Furthermore, the value $\lambda=1.29\times 10^{-3}$ yields the best linear fit, with a slope $\mathbf{M}=0.99\pm 0.26$. The slope is very close to unity, even closer to $\mathbf{M}=1$ than the best-fit value obtained from the comparison between the hydrostatic mass and the observed baryonic mass. Moreover, the value $\mathbf{M}=1$ lies within the uncertainty range and is also located very close to the peak of the corresponding relative likelihood curve, as we can see in Figs. 2(b), (d), (f).
Table 6: Comparison of the goodness-of-fit for different values of the Rastall parameter $\lambda$.
$\lambda$ $\chi^{2}$ $\chi_{\nu}^{2}$ observed mass
0 33184.60 3318.46 baryonic mass
$9.55\times 10^{-2}$ 22.51 2.50 baryonic mass
$1.14\times 10^{-1}$ 55.45 6.16 baryonic mass
0 11.33 1.26 lensing mass
$1.14\times 10^{-3}$ 15.35 1.92 lensing mass
$1.29\times 10^{-3}$ 16.09 2.01 lensing mass
Table 7: Comparison of the goodness-of-fit for Rastall gravity and non-local gravity.
modified gravity $\chi^{2}$ $\chi_{\nu}^{2}$
Rastall gravity, $\lambda=9.55\times 10^{-2}$ 22.51 2.50
Rastall gravity, $\lambda=1.14\times 10^{-1}$ 55.45 6.16
non-local gravity 15.62 1.74
@@
This result indicates that, in the context of the linear fit, Rastall gravity is able to significantly alleviate the hydrostatic mass bias. Nevertheless, a test with a larger sample of galaxy clusters and more lensing mass data is necessary, both to further assess the viability of Rastall gravity and to obtain the more reliable constraint of the value of the Rastall parameter $\lambda$ using observational data. Another noteworthy outcome of this study is that it addresses the issue encountered in Ref. [27], where the hydrostatic mass bias could not be alleviated within the considered framework.
Since there exist uncertainties in the measurements, as mentioned before, it is necessary to evaluate the goodness-of-fit using $\chi^{2}$ and $\chi^{2}_{\nu}$ based on Eqs. (37) and (38), respectively. It is worth noting that larger values of $\chi^{2}$ and $\chi^{2}_{\nu}$ indicate that the measurements deviate more strongly from the values predicted by the model. Moreover, the denominator in the $\chi^{2}_{\nu}$ evaluation contains the number of parameters, which is relevant for Rastall gravity as it introduces an additional degree of freedom, namely $\lambda$.
The results of the goodness-of-fit calculation are presented in Table 6. It can be seen that the improvement in the linear fit (i.e., the slope approaching unity) does not always correspond to a better goodness-of-fit. In particular, although certain values of the Rastall parameter yield slopes that are closer to the expected value, the corresponding $\chi_{\nu}^{2}$ values indicate that the overall agreement with the data is not necessarily improved.
This behavior suggests that, while Rastall gravity is able to capture the global trend of the data (as reflected in the slope), there remain non-negligible discrepancies at the level of individual data points. In other words, the model can improve the overall scaling relation, but does not fully account for the scatter in the observational data.
A similar feature is observed in the second scenario, where the comparison with lensing mass shows that the inclusion of Rastall gravity does not always lead to a better goodness-of-fit compared to the standard (non-Rastall) case. This indicates that, although Rastall gravity can alleviate the hydrostatic mass bias in terms of the slope, it does not systematically improve the statistical consistency with the data.
Note that, in the context of the first scenario where we confront the Rastall version of the hydrostatic mass of galaxy clusters with the observed baryonic mass, we use the dataset of the observed baryonic mass employed in Ref. [33]. In that work, non-local gravity is used as the modified theory of gravity, and the authors obtain the best-fit value of $\mathbf{M}=0.84\pm 0.04$, which is lower than our best-fit value in the first scenario, i.e., $\mathbf{M}=1.07\pm 0.11$. However, in their work, the goodness-of-fit was not evaluated. This motivates us to calculate their goodness-of-fit and compare it with our results.
As shown in Table LABEL:tablechicompare, the comparison with non-local gravity reveals that the latter provides a better goodness-of-fit, despite the fact that the linear relation obtained in the Rastall framework is closer to the expected one. This highlights a trade-off between capturing the global behavior of the data and achieving a statistically better fit.
Overall, these results indicate that Rastall gravity provides a viable phenomenological framework that can improve certain aspects of the mass discrepancy problem, particularly at the level of scaling relations. However, it does not universally outperform other modified gravity models, such as non-local gravity, when evaluated using standard goodness-of-fit criteria. This suggests that further investigations, including larger data samples and more refined modeling, are necessary to fully assess its effectiveness.
@@ 4 Conclusion
In this study, we have derived the formalism of the hydrostatic mass of galaxy clusters within the framework of Rastall gravity. We find that Rastall gravity has a nontrivial influence on the hydrostatic mass of galaxy clusters. Two scenarios are considered: (i) the absence of DM, in which the Rastall hydrostatic mass is compared with the observed baryonic mass; and (ii) the existence of DM, where the Rastall hydrostatic mass is employed to alleviate the hydrostatic mass bias.
@@
On the other hand, in the second scenario, by using nine galaxy clusters as the sample, we find that the best-fit value, $\mathbf{M}=0.99\pm 0.26$, is obtained for $\lambda=1.29\times 10^{-3}$. In this case, unity lies very close to the peak of the corresponding relative likelihood curve in the comparison between the Rastall hydrostatic mass and the observed lensing mass. In the context of the scaling relation, this result indicates that Rastall gravity can effectively alleviate the hydrostatic mass bias problem.
We also calculate the goodness-of-fit for each case, using chi-squared $\chi^{2}$ and reduced chi-squared $\chi^{2}_{\nu}$ evaluations. The combination of the linear fit and the goodness-of-fit evaluations indicates that Rastall gravity provides a phenomenological framework that can improve certain aspects of the mass discrepancy problem, particularly at the level of scaling relations. However, it does not universally outperform other modified gravity models, such as non-local gravity, when evaluated using standard goodness-of-fit criteria.
Nevertheless, a larger sample of galaxy clusters in the comparison between the Rastall hydrostatic mass and the gravitational lensing mass is required to further assess the viability of Rastall gravity and to constrain the Rastall parameter $\lambda$ using observational data.
@@
During the preparation of this work, M. Lawrence Pattersons used ChatGPT (OpenAI) in order to support brainstorming and organization of some theoretical ideas, and to improve the language and readability of the manuscript. All physical interpretations, model constructions, calculations, and final manuscript preparation were performed by the authors, who take full responsibility for the content. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the paper.
@@ Acknowledgements
Appendix A Statistical data of the linear fits
The statistical data of the linear fits are shown by Table 8.